### v2

# A Proof of the Weak Gravity Conjecture via Leading-Order Quantum Gravity Effects

Abstract
The weak gravity conjecture posits that there should exist at least one state with a charge-to-mass ratio $q/M>1$, ensuring that an extremal charged black hole can still decay. In this paper, we study the corrections to the thermodynamics of Reissner-Nordström black holes within an effective field theory framework. The effective action originates from integrating out massless particles, including gravitons, at one-loop level. We demonstrate that these quantum gravity effects naturally give rise to the super-extremality of charged black holes, which aligns with a recent result by Arkani-Hamed et al. Furthermore, our formalism provides a comprehensive thermodynamic analysis of both extremal and non-extremal RN black holes, offering additional interesting implications.
I Introduction
In Einstein’s general relativity, the extremality bound for a Reissner-Nordström (RN) black hole is given by $\frac{q}{M}\leq 1$ in natural units, where $q$ (implicitly $|q|$) and $M$ are the electric charge and mass of the black hole, respectively. For an extremal charged black hole, if it is assumed to decay into two or more lighter components, then, due to the conservation of energy and charge, one of the decay products must have a charge-to-mass ratio $\frac{q}{M}>1$. If such a super-extremal object does not exist in nature, the decay process would be completely forbidden, and the evolution of black holes would cease at a final stage. Eventually, the universe would be populated by a large number of stable remnants with various charges, which is an unfavorable scenario.
This situation has motivated the so-called Weak Gravity Conjecture (WGC). The WGC suggests that for any $U(1)$ gauge field coupled to gravity, there must exist at least one object with charge $q$ and mass $M$ such that its charge-to-mass ratio exceeds the bound of a large extremal black hole Arkani-Hamed:2006emk . As the name implies, consistent theories must include states where electromagnetic repulsion dominates gravitational attraction. Currently, there are various versions of the WGC, both mild and strong, with different motivations, interpretations, and implications. For an in-depth review of this field, see Ref. Harlow:2022ich .
Proving the WGC is a challenging task with multiple approaches. A natural idea is that the well-known bound $q\leq M$ is a result of Einstein’s gravity, which is a classical theory from a modern perspective. Thus, when quantum gravity effects are properly considered, the black hole may possess a state with $q>M$, allowing an extremal charged black hole to decay further.
In this regard, higher-derivative terms should be incorporated into the gravitational effective action as manifestations of quantum gravity effects Boulware:1985wk ; Cardoso:2018ptl ; BarrosoVarela:2023ull . The quantum-corrected effective action can be formally written as
$\begin{split}I_{\text{tot}}&=\int_{\mathcal{M}}\frac{1}{16\pi}(R-F_{\mu\nu}F^{\mu\nu})+\epsilon L_{\text{hd}},\end{split}$ (1)
where $L_{\text{hd}}$ represents the higher-derivative terms and the parameter $\epsilon$ is introduced to track their effects. Researchers can study the variations in the mass and entropy of an extremal RN black hole caused by turning on these terms. Typically, a decrease in the mass $M_{\text{ext}}$ at a fixed charge $q$ is expected. Therefore, proving the WGC is equivalent to verifying whether the following expression holds:
$\left(\frac{\partial M_{\text{ext}}(q,\epsilon)}{\partial\epsilon}\right)_{q}<0.$ (2)
Significant achievements have been gained in recent years Kats:2006xp ; Cremonini:2009ih ; Cremonini:2019wdk ; Cheung:2018cwt ; Cheung:2019cwi ; Noumi:2022ybv . Specifically, four-derivative terms such as $R^{2}$, $R_{\mu\nu}^{2}$ and $R_{\mu\nu\alpha\beta}^{2}$ have been extensively studied, and eq.(2) is utilized to place constraints on the combination of the coefficients of these curvature terms. There is a growing body of evidence in support of the WGC, coming from concrete models constructed from string theories or under certain assumptions. In these studies, the local $\mathcal{O}(R^{2})$ terms are often regarded as making the main contributions to the relevant corrections. However, the loop effects of massless particles including gravitons can dominate over them in the IR and provide the leading contributions Harlow:2022ich .
Recently, the loop effects of massless particles have been studied by Arkani-Hamed et al. in Ref. Arkani-Hamed:2021ajd . In this paper, we will start from a non-local effective action that captures these loop effects and demonstrate that such effects indeed reduce the mass of an extremal RN black hole. This result confirms the findings of Arkani-Hamed:2021ajd , offering a fresh and concise perspective on the problem. Moreover, we present a detailed analysis of the thermodynamics of both extremal and non-extremal RN black holes. We emphasize that the results are model-independent, as they depend solely on the low-energy particle spectrum of nature and are insensitive to the UV details of quantum gravity theories.
II One-Loop Effective Action
Thanks to the pioneering works Donoghue:1994dn ; Donoghue:2014yha ; Barvinsky:1985an ; Calmet:2019eof ; Calmet:2024neu ; Teixeira:2020kew , after integrating out the quantum fluctuations of gravitons and other massless particles at one-loop level, the higher-derivative term is given by
$\begin{split}&L_{\text{hd}}=\left[c_{1}(\mu)R^{2}+c_{2}(\mu)R_{\mu\nu}R^{\mu\nu}+c_{3}(\mu)R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}\right]\\ &\!-\!\left[\alpha R\ln(\frac{\Box}{\mu^{2}})R\!+\!\beta R_{\mu\nu}\ln(\frac{\Box}{\mu^{2}})R^{\mu\nu}\!+\!\gamma R_{\mu\nu\alpha\beta}\ln(\frac{\Box}{\mu^{2}})R^{\mu\nu\alpha\beta}\right]\!.\end{split}$ (3)
Note that the loops involving massless particles generate nonlocal effects, which are reflected by the presence of the nonlocal operator $\ln\left(\frac{\Box}{\mu^{2}}\right)$, where $\Box$ is the d’Alembert operator. By construction, the effective action is renormalization group invariant, that is, it is independent of the renormalization parameter $\mu$ El-Menoufi:2015cqw . For the convenience of manipulation, we will absorb $c_{i}(\mu)$ into its partner $\ln\left(\frac{\Box}{\mu^{2}}\right)$, thereby defining the operator $\ln\left(\frac{\Box}{\mu_{*}^{2}}\right)$. Here, $\mu_{*}$ represents the characteristic energy scale beyond which the effective field-theoretical description becomes invalid. The value of $\mu_{*}$ should be determined by the UV theory of quantum gravity Calmet:2024neu . Without loss of generality, it can be taken to be near the Planck energy. It should be noted that we always work in a regime $r>1/\mu_{*}$, which is below such an energy scale.
We are particularly interested in this effective action for two main reasons. First, the Wilson coefficients $\alpha$, $\beta$, and $\gamma$ of the nonlocal terms are calculable and insensitive to concrete UV theories. This insensitivity represents a model-independent prediction of the effective theory. For the convenience of the reader, we list these coefficients below Donoghue:2014yha ; qfcs :
$\alpha$ $=\frac{1}{11520\pi^{2}}(5(6\xi\!-\!1)^{2}N_{s}\!-\!5N_{f}\!-\!50N_{v}\!+\!430N_{g}),$ (4)
$\beta$ $=\frac{1}{11520\pi^{2}}(-2N_{s}+8N_{f}+176N_{v}-1444N_{g}),$ (5)
$\gamma$ $=\frac{1}{11520\pi^{2}}(2N_{s}+7N_{f}-26N_{v}+424N_{g}),$ (6)
where $N_{s}$, $N_{f}$, $N_{v}$, and $N_{g}$ denote the numbers of scalars, four-component fermions, vectors, and gravitons in the low-energy particle spectrum of nature, respectively.
Second, the nonlocal terms are at the leading order in the curvature expansions, which makes their effects the most prominent. Specifically, even if one takes into account the additional local curvature terms $L_{\text{hd}}=c_{2}R^{2}+c_{3}R^{3}+c_{4}R^{4}+\cdots$ that are constructed from a UV theory in a top-down manner, their contributions are overshadowed by those of the $R\log\Box R$ terms. According to the scaling behaviors of these terms, their contributions can be arranged in a sequence: $\log(r^{2}\mu_{*}^{2})$, $1$, $\frac{1}{r^{2}}$, $\frac{1}{r^{4}}$, $\cdots$. When working in the regime $r>1/\mu_{*}$, the first term dominates over all other contributions. In contrast, when approaching the scale $r\sim 1/\mu_{*}$, all contributions are expected to be of the same order, and the effective field theory becomes invalid. Only a complete quantum gravity theory can sum over all these contributions. Since we are working within the regime $r>1/\mu_{*}$, we will focus on the nonlocal terms and analyze their implications for the WGC, while ignoring all the contributions from local terms.
Besides, the higher curvature terms (3) shares some similarities with that analyzed in Ref.Arkani-Hamed:2021ajd through field redefinitions, and it may be equivalent in some deeper sense, as both ultimately describe the same physics of loop effects. However, the technical challenges faced in analyzing eq.(3) are different. The primary difficulty here lies in how to handle the non-local operator $\ln\frac{\Box}{\mu_{*}^{2}}$ and its variation; see, e.g., Donoghue:2014yha ; Calmet:2019eof ; Xiao:2021zly .
III Thermodynamic Corrections of the RN Black Hole
To prove eq.(2), we need to calculate the corrections to the thermodynamic quantities of a RN black hole, such as its mass and entropy, which are induced by switching on $\epsilon L_{\text{hd}}$.
Conventionally, the expressions of thermodynamic quantities should be derived from the black hole metric using standard techniques. At the first order in $\epsilon$, the perturbed RN black hole metric can be formally written as $g_{\mu\nu}=\bar{g}_{\mu\nu}+\epsilon g_{\mu\nu}^{(1)}$, where $\bar{g}_{\mu\nu}$ represents the original unperturbed RN metric $ds^{2}=-f(r)dt^{2}+f^{-1}(r)dr^{2}+r^{2}d\Omega^{2}$, and $f(r)=1-\frac{2M}{r}+\frac{q^{2}}{r^{2}}$. However, determining the exact form of the perturbed metric induced by the nonlocal terms in eq.(3) is far from an easy task. Due to the lack of appropriate techniques to handle the variation $\delta\ln\Box$, the gravitational field equations cannot be derived straightforwardly from the variation analysis of the Lagrangian. Thus, certain assumptions have to be made to obtain the perturbed field equations and the corresponding black hole solution. This has been done in our previous work Xiao:2021zly .
Fortunately, there exists a convenient approach indicating that, at the first order in $\epsilon$, the corrections to the black hole’s thermodynamic quantities from any higher-derivative terms $\epsilon L_{\text{hd}}$ can be directly obtained using the unperturbed metric. This approach was discovered occasionally by Guber et al. Gubser:1998nz and rediscovered and well-explained recently by Reall and Santos Reall:2019sah . For more discussions on its validity and applications, refer to Refs.Caldarelli:1999ar ; Xiao:2022auy ; Xiao:2023two ; Ma:2023qqj . By following this approach, we can deduce the expression of $M_{ext}(q,\epsilon)$ without the need of making unnecessary assumptions.
We now start to evaluate the onshell Euclidean action $I_{\text{tot}}(g_{\mu\nu})=I_{\text{RN}}+\epsilon I_{\text{hd}}$ by substituting $g_{\mu\nu}$ into it. The Euclidean action is directly related to the partition function and free energy of the system. According to the aforementioned approach, when evaluating the Euclidean action at the first order in $\epsilon$, using the perturbed metric $g_{\mu\nu}$ or the unperturbed metric $\bar{g}_{\mu\nu}$ can yield the same result. The reasoning is as follows. The Euclidean action consists of two parts, $I_{\text{RN}}$ and $\epsilon I_{\text{hd}}$, which can be evaluated separately. First, consider the two-derivative part $I_{\text{RN}}\equiv\frac{1}{16\pi}\int_{M}(R-F_{\mu\nu}F^{\mu\nu})$ along with its boundary terms. Since the original RN metric $\bar{g}_{\mu\nu}$ is the extremum of this functional integral, any perturbation around $\bar{g}_{\mu\nu}$ will only cause a correction of $\mathcal{O}(\epsilon^{2})$. That is, $I_{\text{RN}}(g_{\mu\nu})=I_{\text{RN}}(\bar{g}_{\mu\nu})+\mathcal{O}(\epsilon^{2})$ under the condition that $g_{\mu\nu}$ and $\bar{g}_{\mu\nu}$ have the same boundary conditions  [^1] It is understandable that, for a reasonable comparison between $\mathcal{F}[f_{1}(x)]$ and $\mathcal{F}[f_{2}(x)]$ in functional analysis, the two functions $f_{1}(x)$ and $f_{2}(x)$ must satisfy the same boundary conditions.. This analysis also holds when changing the thermodynamic ensemble Cremonini:2019wdk . As a result, when evaluating $I_{\text{RN}}(g_{\mu\nu})$ in an ensemble depending on temperature and charge, the corresponding free energy can be written as
$F_{RN}(T,q)=\frac{3q^{2}+\bar{r}_{h}^{2}}{4\bar{r}_{h}}+\mathcal{O}(\epsilon^{2}).$ (7)
The first term is directly taken from the known formula for the unperturbed RN black hole with the same temperature $T$ and charge $q$. So $\bar{r}_{h}$ represents the outer horizon radius of the unperturbed black hole, and the expression of $\bar{r}(T,q)$ should be solved from the relation
$T=\frac{1}{4\pi\bar{r}_{h}}-\frac{q^{2}}{4\pi\bar{r}_{h}^{3}},$ (8)
and then inserted into eq.(7). In practice, this is not necessary because we can use the chain rule to compute quantities such as $\frac{\partial F}{\partial T}$. Technically, we could also re-express eq.(7) as a function of the horizon radius $r_{h}=\bar{r}_{h}+\mathcal{O}(\epsilon)$ of the perturbed black hole, but it is more convenient to use $\bar{r}_{h}$. For example, taking the external limit $T\rightarrow 0$ is equivalent to taking $\bar{r}_{h}\rightarrow q$. If we use $r_{h}$ as the variable, all such expressions will soon become complicated. For those who are confused in thinking back and forth between the characteristic quantities of a perturbed and unperturbed black hole, it is helpful to ignore the origin of $\bar{r}_{h}$ and simply regard it as a convenient parameter related to the temperature $T$ and charge $q$ through eq.(8). Then, one can focus on the perturbed black hole that we are actually concerned with.
Next, consider the second part $\epsilon I_{\text{hd}}\equiv\epsilon\int_{\mathcal{M}}L_{\text{hd}}$. It is obvious that $\epsilon I_{\text{hd}}(g_{\mu\nu})=\epsilon I_{\text{hd}}(\bar{g}_{\mu\nu})+\mathcal{O}(\epsilon^{2})$ due to the presence of the factor $\epsilon$. In evaluating $\epsilon I_{\text{hd}}(\bar{g}_{\mu\nu})$, it would be tricky to calculate the terms like $\ln(\frac{\Box}{\mu_{*}^{2}})R_{\mu\nu\alpha\beta}$. However, from dimensional analysis, we can anticipate that its main contribution has the form $-\ln(r^{2}\mu_{*}^{2})R_{\mu\nu\alpha\beta}$, and the explicit calculation can be found in, e.g., Calmet:2019eof ; Delgado:2022pcc ; El-Menoufi:2017kew . Substituting it into $\epsilon L_{hd}$ and calculating the Euclidean integral from the horizon to infinity, we can obtain the expression of $\epsilon I_{\text{hd}}(\bar{g}_{\mu\nu})$.
In total, at the first order in $\epsilon$, the free energy of the perturbed RN black hole is given by
Since the RN black hole is Ricci-flat, the Wilson coefficient $\alpha$ does not contribute here. To highlight the universal logarithmic corrections, we have omitted the subleading terms, as $\log(\bar{r}_{h}^{2}\mu_{*}^{2})\gg 1$. These subleading terms are also non-universal because their values can be easily altered by other model-dependent local curvature-squared interactions.
All thermodynamic quantities can be derived from the free energy  [^2] The computations of this study are available in the auxiliary file of the arXiv version of the paper.. The result is consistent with our previous work Xiao:2021zly , where we explicitly solved the perturbed black hole metric. Specifically, the expressions for the mass and entropy are
and
IV Extermal RN black hole and the WGC
Finally, we are in a position to directly examine the WGC. By taking the extremal limit $T\rightarrow 0$, which is equivalent to $\bar{r}_{h}\rightarrow q$, we get
and
In these equations, we have substituted the values of the Wilson coefficients $\alpha$, $\beta$, and $\gamma$. As a consistency check, eq.(13) reproduces the logarithmic correction to black hole entropy obtained from an independent approach by Sen in Sen:2012kpz , except for the coefficient of gravitons. Unlike other particles, the coefficient of gravitons may depend on the gauge choice and the subtleties associated with dealing with zero modes Calmet:2024neu ; Kallosh:1978wt .
From the expression of $M_{ext}$, we can straightforwardly obtain the core result
$\left.\frac{dM_{\text{ext}}}{d\epsilon}\right|_{q}=-\frac{\mathcal{N}}{600\pi r_{h}}\log(\bar{r}_{h}^{2}\mu_{*}^{2}),$ (14)
where
$\mathcal{N}\equiv n_{s}+6n_{f}+12n_{v}+42n_{g}.$ (15)
Interestingly, the coefficients conspire to produce a definite sign. Since $n_{s}$, $n_{f}$, $n_{v}$, and $n_{g}$ represent the numbers of particles in the spectrum, it is certain that $\mathcal{N}>0$. Consequently, quantum gravity effects indeed reduce the mass of an extremal charged black hole at a fixed charge $q$, as first pointed out in Ref. Arkani-Hamed:2021ajd .
As mentioned above, all coefficients are fixed except for the ambiguity associated with the coefficients of gravitons. We have also tested graviton coefficients from Calmet:2024neu , which are claimed to be gauge-invariant and unique. In this case, the graviton coefficient in eq.(15) becomes $242$ instead of $42$, but the sign remains definite.
The characteristic energy scale $\mu_{*}$ is determined by the UV theory of quantum gravity. Thus, the presence of $\ln\mu_{*}$ in eq.(14) indicates the UV sensitivity of our result. However, the model-dependent value of $\mu_{*}$ does not affect our conclusion for the WGC because we are working within the regime $r>1/\mu_{*}$, where the effective field-theoretical description is valid. The physics near the scale $r\sim 1/\mu_{*}$ would be described by a complete theory of quantum gravity and is not relevant to our current analysis.
Computing $\frac{\partial M_{ext}}{\partial\epsilon}$ is straightforward. However, we can also derive eq.(14) from other formulas, which offer certain advantages. There is a crucial formula, proposed and proven in Goon:2019faz , given by
$\left(\frac{\partial M_{\text{ext}}(q,\epsilon)}{\partial\epsilon}\right)_{q}\!=\!-\!\lim\limits_{M\rightarrow M_{\text{ext}}}T\left(\frac{\partial S(M,q,\epsilon)}{\partial\epsilon}\right)_{M,q}.$ (16)
In fact, from the extended first law $dM=TdS+\Phi dq+U_{\epsilon}d\epsilon$, it can be easily observed that both sides of eq.(16) are equal to $\lim\limits_{T\rightarrow 0}U_{\epsilon}=\lim\limits_{T\rightarrow 0}\frac{\partial F}{\partial\epsilon}$ Xiao:2023two . Therefore, $\frac{\partial M_{ext}}{\partial\epsilon}$ can be evaluated as the spatial integral $-\int d^{3}x\sqrt{g}L_{\text{hd}}$ for an extremal RN black hole Cheung:2018cwt ; Cheung:2019cwi . This provides a rapid way to check whether the WGC is satisfied for a given higher-derivative term $\epsilon L_{\text{hd}}$ without delving into the analysis of the perturbed black hole thermodynamics.
Another way is to evaluate the shift in entropy due to switching on the quantum gravity effects. The formula (16) is significant as it links the mass decrease required by the WGC to the entropy increase. Notably, our analysis further establishes a novel connection between the required super-extremality and the logarithmic entropy corrections, that is, both of them are generated by the same quantum gravity effects. Notice that there is a caveat here. The derivative $T\left(\frac{\partial S}{\partial\epsilon}\right)_{M,q}$ is evaluated at a fixed $M$ rather than at a fixed $T$. Since turning on $\epsilon$ while holding $M$ constant will shift the temperature away from its extremal value Harlow:2022ich , we should use eq.(11) rather than the more familiar eq.(13) to derive the results. This explains why no clues regarding the WGC have previously been derived from the logarithmic entropy corrections, despite their existence in the literature for many years.
The WGC is mainly concerned with the fate of extremal RN black holes. However, we can also explore the quantum gravity effects on non-extremal RN black holes to see how the charge-to-mass ratio changes Solomon:2020pja . In this regard, we find that
$\left.\frac{\partial M}{\partial\epsilon}\right|_{T,q}=-\frac{\mathcal{N}}{300\pi}\frac{q^{4}\left(2\bar{r}_{h}^{2}-q^{2}\right)\log\left(\bar{r}_{h}^{2}\mu_{*}^{2}\right)}{\bar{r}_{h}^{5}\left(\bar{r}_{h}^{2}-3q^{2}\right)}.$ (17)
The singularity at $\bar{r}_{h}=\sqrt{3}q$ indicates a critical point that coincides with the boundary between the thermodynamically stable and unstable RN black holes, as determined by the sign of the heat capacity $C_{q}$. Thus, for a thermodynamically stable black hole, the value of $q/M$ increases, while for an unstable one, it decreases. A general discussion of the correlation between thermodynamic stability and super-extremality can be found in Cheung:2018cwt , where it was first formalized based on tree-level higher derivative terms.
V Concluding remarks
In this paper, we have explored the effective action obtained by integrating out massless particles at the one-loop level. We confirm that the loop effects of quantum gravity can naturally induce super-extremality ($q>M$) in charged black holes Arkani-Hamed:2021ajd . It suggests that quantum gravity plays a crucial role in facilitating the decay of extremal black holes into more fundamental components. These effects thus have far-reaching implications for our understanding of the ultimate evolution of black holes and the universe.
We emphasize that these results are independent of specific UV theories of quantum gravity. Instead, they depend solely on the low-energy particle spectrum of nature, which endows them with predictive power. Specifically, by using the particle spectrum of the Standard Model with a minimally coupled Higgs and taking $\mu_{*}$ as the Planck energy, we predict that for a black hole with a charge of approximately $10^{-16}C$ (the typical charge carried by a dust particle), the quantity $\frac{q}{M}-1\approx 5.2\times 10^{-5}$.
Finally, we have provided a comprehensive thermodynamic analysis of both extremal and non-extremal RN black holes, which has further interesting implications. For extremal RN black holes, it unexpectedly reveals a connection between super-extremality and the logarithmic entropy corrections, which were previously viewed as two independent research directions. For non-extremal RN black holes, we also demonstrate a relationship between the shift in the charge-to-mass ratio and the thermodynamic stability properties of the black holes. Both of these aspects merit further investigation.
Acknowledgments
We thank Grant Remmen, Nima Arkani-Hamed, and Yu-Tin Huang for informing us of their earlier work on loop effects of massless particles and for the constructive communications after our work was posted on arXiv. XY would like to thank Yu Tian and Hongbao Zhang for useful discussions. YX is also thankful to the Higgs Centre for Theoretical Physics at the University of Edinburgh for providing research facilities and hospitality during the visit. This work was supported in part by China Scholarship Council with Grant No. 202408130101, the Hebei Natural Science Foundation with Grant No. A2024201012, and the National Natural Science Foundation of China with Grant No. 12475048.